A hockey ball is struck by a hockey stick. The mass of the ball (62g), the mass of the head of the hockey stick (206g), the velocity of the head of the hockey stick immediately prior to impact (34m/s) and the coefficient of restitution (0.9). Using these values calculate how fast the ball is moving immediately after impact? Round your answer to one decimal place


1
Expert's answer
2021-05-03T07:09:26-0400
"\frac{u+v'}{v}=0.9\\Mv=-Mv'+mu\\Mv=-M(0.9v-u)+mu
\\206(34)=-206(0.9(34)-u)+62u\\u=49.7\frac{m}{s}"